A pendulum takes 30 sec to complete 20 oscillition . What is the time period ?
 

Grade:7

3 Answers

Venkat
5 years ago
Given that the pendulum makes 20 oscillation in 30secs, the oscillation made in one sec is,oscillations = 20/30 = 2/3 = 0.6667
Venkat
5 years ago
Given that the pendulum makes 20 oscillation in 30secs, the oscillation made in one sec is,
 
oscillations made per second = 20/30 = 2/3 = 0.6667.
 
Recall that the number of oscillations made per second is called as frequency. Hence,
Frequency = 0.6667 Hz
Time period = 1/ Frequency.
Time period = 1/0.6667 = 1.5 sec.
 
Hence, time period is 1.5 sec
 
Scholr Boy
5 years ago
Use Scholr App for Exam Preparation.
Given that the pendulum makes 20 oscillation in 30secs, the oscillation made in one sec is,
 
oscillations made per second = 20/30 = 2/3 = 0.6667.
 
Recall that the number of oscillations made per second is called as frequency. Hence,
Frequency = 0.6667 Hz
Time period = 1/ Frequency.
Time period = 1/0.6667 = 1.5 sec.
 
Hence, time period is 1.5 sec